OpenStudy (anonymous):

Stumped by rational expressions, please walk me through the steps if possible.

OpenStudy (anonymous):

Question is: \[x^2-16/2x^2-9x+4\div2x^2+14+24/4x+4\]

OpenStudy (luigi0210):

Uhm, is it \[\LARGE \frac{x^2-16}{2x^2-9x+4} \div \frac{2x^2+14x+24}{4x+2}\] ?

OpenStudy (anonymous):

Yes, sorry I didn't know how to make fractions with the equation tool. I also didn't know how to put equations my original question.

OpenStudy (luigi0210):

Whoops, typo, 2, not 4, in the bottom second equation right? But when you are dividing fractions, switch the sign ti multiplication and flip the second fraction.

OpenStudy (luigi0210):

I mean 4, not 2

OpenStudy (anonymous):

Yeah, it's a 4, so it would be \[\frac{ x^2-16 }{ 2x^2-9x+4 }\times \frac{ 4x+4 }{2x^2 }\]?

OpenStudy (anonymous):

... it got cut off?

OpenStudy (luigi0210):

I get what you mean, but yea, now, factor everything.

OpenStudy (anonymous):

I'm a little slow at this it'll probably take me a minute or two

OpenStudy (luigi0210):

It's alright, take you time.

OpenStudy (anonymous):

\[\frac{ (x-4)(x+4) }{ (2x-1)(x-4) }\]

OpenStudy (anonymous):

That's the first half more coming, sorry I'm doing the work on paper, then trying to copy it over to this >_<

OpenStudy (anonymous):

\[\frac{ 2(1+2x) }{ 2(19+x^2) }\]

OpenStudy (luigi0210):

I think you might need to re-check that. Here's what it should be: \(\LARGE \frac{x^2-16}{2x^2-9x+4} \div \frac{2x^2+14x+24}{4x+4}==>\frac{(x+4)(x-4)}{(2x-1)(x-4)}*\frac{4(x+1)}{2(x+3)(x+4)}\)

OpenStudy (luigi0210):

\(\LARGE \frac{x^2-16}{2x^2-9x+4} \div \frac{2x^2+14x+24}{4x+4}=>\frac{(x+4)(x-4)}{(2x-1)(x-4)}*\frac{4(x+1)}{2(x+3)(x+4)}\)

OpenStudy (anonymous):

So, I was right for the first half... Then no clue what happened to be honest

OpenStudy (luigi0210):

Yup, now all we do is just start canceling what we can, do you see anything we can cancel?

OpenStudy (anonymous):

x-4,x+4...

OpenStudy (luigi0210):

Right! So those cancel out  So those cancel.. there is one more thing we can simplify, do you know what that is?